OpenStudy (anonymous):

D22: One line passes through the points (−12,16) and (−1,38). A second line has the equation 2y+x=23. Are these lines parallel, perpendicular, or neither? Discuss your solution strategy and then implement it to demonstrate your answer.

OpenStudy (texaschic101):

compare the slopes 2y + x = 23 -- find the slope by putting it in y = mx + b form, and m is your slope. 2y = -x + 23 y = -1/2x + 23/2 slope is -1/2 given 2 points, we can use the slope formula to find the slope slope = (y2 - y1) / (x2 - x1) (-12.16) x1 = -12 and y1 = 16 (-1,38) x2 = -1 and y2 = 38 now we sub slope = (38 - 16) / (-1 - (-12) slope = 22/(-1 + 12) slope = 22/11 slope = 2 so our slopes are -1/2 and 2 parallel lines will have the same slope. perpendicular lines will have negative reciprocal slopes. All that means is " flip " the slope and change the sign. For example, you have a slope of 4, which is the same as 4/1, if you " flip " the slope and change the sign, -1/4, it is perpendicular. so what do you think ?